# Bringing memory to Boolean networks: a unifying framework

### v1

@@ 2.7 Cuttable updates
The asynchronous graph of $\mathsf{C}$ is given here below with an example of a $\mathtt{C}$-trajectory. The same trajectory is also given in Figure 2 in a different format.
000000000000001001001001010010010010011011011011100100100100101101101101110110110110111111111111
@@ 00 $\begin{pmatrix}0&0&0\\ 0&0&0\\ 0&0&0\end{pmatrix}$ $000$
$1$ $1$ $\begin{pmatrix}0&0&0\\ 0&0&0\\ 0&0&0\end{pmatrix}$ $000$ $000$ $100$
$2$ $2$ $\begin{pmatrix}0&0&0\\ 1&0&0\\ 0&0&0\end{pmatrix}$ $100$ $100$ $110$
$3$ $3$ $\begin{pmatrix}0&0&0\\ 1&0&0\\ 0&2&0\end{pmatrix}$ $010$ $110$ $111$
@@ 3.1 Hierarchy by reachability and trajectory for the different update modes
We are interested in comparing the collections of reachability pairs amongst the different update modes listed above. For any update mode $\mu\in\{\mathtt{A},\mathtt{H},\mathtt{T},\mathtt{M},\mathtt{S},\mathtt{I},\mathtt{C}\}$, we denote its collection of reachability pairs by $\mu^{*}=\{(x,y):x\to_{\mu}^{*}y\}$. Then $\nu$ weakly simulates $\mu$ if $\mu^{*}\subseteq\nu^{*}$ for all BNs. In Figure 1, a strict containment $\mu^{*}\subset\nu^{*}$ means that the containment $\mu^{*}\subseteq\nu^{*}$ holds for all BNs and that there exists a BN where this containment is strict.
Our hierarchy by reachability is based on a hierarchy by trajectories. As such, we also give the full hierarchy of trajectories in Figure 1. Again, a strict containment $\mu\subset\nu$ means that the containment $\mu\subseteq\nu$ holds for all BNs and that there exists a BN where this containment is strict. Clearly, if $\mu\subseteq\nu$, then $\mu^{*}\subseteq\nu^{*}$.
@@
Figure 1: Hierarchy of different update modes.
Figure 2: Possible trajectories of certain BNs according to different update modes.
The rest of this subsection is devoted to the proof of Theorem 3.1, broken down into items a to h.
@@
    $\mathtt{I}\not\subseteq\mathtt{T}$.
Example 3.2 (An interval trajectory that is not trapping).
We have already shown that it is an $\mathtt{I}$-trajectory. It is not a $\mathtt{T}$-trajectory, as $x_{1}^{4}\neq x^{3}_{1}$ implies that $i^{4}=1$, while $\mathsf{I}_{1}(x^{0})=\mathsf{I}_{1}(x^{1})=\mathsf{I}_{1}(x^{2})=\mathsf{I}_{1}(x^{3})=1$ implies that $s^{4}\notin\{x^{0},\dots,x^{3}\}$.
  1. (e)
@@
We now prove that the trapping update mode and the subcube-based update mode allow to reach any configuration in the whole principal trapspace.
@@
Let $x=x^{0}\to_{\mathtt{I}}\dots\to_{\mathtt{I}}x^{l}=y$ be an interval trajectory where $d_{\mathrm{H}}(s^{a-1},s^{a})\leq 1$ for all $11\leq a\leq p-$ 1 and $d_{\mathrm{H}}(s^{p-1},s^{p})>1$. Let us prove how we can replace the transition $x^{p}\to_{\mathtt{I}}x^{p+1}$ with a sequence of transitions $u^{0}=x^{p}\to_{\mathtt{I}}\dots\to_{\mathtt{I}}u^{n}=x^{p+1}$ and still have an interval trajectory.
@@
is an interval trajectory. The first part, $\dot{x}^{0}\to_{\mathtt{I}}\dots\to_{\mathtt{I}}\dot{x}^{p}$, follows from our hypothesis. For the second part, $\dot{x}^{p}\to_{\mathtt{I}}\dots\to_{\mathtt{I}}\dot{x}^{p+n}$, use $w^{p+i}=(i^{p},v^{i-1},u^{i-1})$. For the third part, $\dot{x}^{p+n}\to_{\mathtt{I}}\dots\to_{\mathtt{I}}\dot{x}^{n+1-l}$, then $\dot{w}^{n-1+a}=w^{a}$ is a valid triple for interval updates. ∎
@@
Let $x=x^{0}\to_{\mathtt{I}}x^{1}\to_{\mathtt{I}}\cdots\to_{\mathtt{I}}x^{l}=y$ be an interval trajectory with corresponding triples $w^{a}=(i^{a},s^{a},t^{a})$ for all $1\leq a\leq l$.
According to Lemma 3.5, we assume $d_{\mathrm{H}}(s^{a-1},s^{a})\leq 1$ for all $1\leq a\leq l$.
We first prove, by induction over $1\leq a\leq l$, that there always exists a history-based trajectory $x=x^{0}\to_{\mathtt{H}}s^{1}\to_{\mathtt{H}}\cdots\to_{\mathtt{H}}s^{a}$. The basic case $a=1$ is trivial because $s^{1}=x^{0}$. Assume the statement holds for $a-1$, i.e. $x=x^{0}\to_{\mathtt{H}}s^{1}\to_{\mathtt{H}}\cdots\to_{\mathtt{H}}s^{a-1}$ is a history-based trajectory. Lemma 3.4 settles the case where $s^{a}=s^{a-1}$, hence we assume $s^{a-1}\neq s^{a}$ and, according to Lemma 3.5, $\Delta(s^{a-1},s^{a})=\{j\}$. Then, for $q=V^{a}_{j}$, we have $q<a$ and $x^{q}_{j}=s^{a}_{j}$. Suppose, for the sake of contradiction, that $x^{m}_{j}=x^{q}_{j}$ for all $0\leq m\leq q$. Then, $s^{m}_{j}=x^{q}_{j}$ for all $0\leq m\leq q$, and in particular $s^{a-1}_{j}=s^{a}_{j}$, which is the desired contradiction. Therefore, there exists $q^{\prime}<q$ such that $f_{j}(s^{q^{\prime}})=x^{q}_{j}$. Thus, let $s^{a}=s^{q^{\prime}}$ and $i^{a}=j$. We obtain $(f_{j}(s^{q^{\prime}})=s^{a}_{j},s^{a-1}_{-j})=s^{a}$.
We now prove that we can extend the history-based trajectory all the way to $y$. More formally, we prove that $x=x^{0}\rightarrow s^{1}\rightarrow\cdots\rightarrow s^{l}\rightarrow\cdots\rightarrow y$ is a history-based trajectory (where $s^{l}\rightarrow\cdots\rightarrow y$ is any geodesic). Let $\Delta=\Delta(s^{l},y).$ For all $j\in\Delta$, there exists $q_{j}<l$ such that $f_{j}(s^{q_{j}})=y_{j}$. Since $s_{j}^{l}=x^{m}_{j}\neq y_{j}$ (for some $m<l$), there exists $q_{j}$ such that $j=i^{q_{j}}$ and $f_{j}(s^{q_{j}})=y_{j}$. Then $x=x^{0}\rightarrow s_{1}\rightarrow\cdots\rightarrow s_{l}\rightarrow g^{l+1}\cdots\rightarrow g^{l+\mid\Delta\mid}=y$ is a history-based trajectory with $\{i^{l+1},\ldots,i^{l+\mid\Delta\mid}\}=\Delta$, and for $l+1\leq m\leq l+\mid\Delta\mid$, let $s^{m}=s^{q_{m}}$ when $f_{i_{m}}(s^{q_{m}})=y_{i_{m}}$ and finally $t^{m}=g^{m-1}$. ∎
@@
All the update modes considered in this paper have the property that trajectories can be compressed. Our definition of trajectory allows for transitions $x^{a-1}\to_{\mu}x^{a}$ where $x^{a-1}=x^{a}$; a simple example is when $x^{a-1}$ is a fixed point of $f$ and $s^{a}=t^{a}=x^{a-1}$. We can compress a trajectory by removing consecutive duplicates and obtain a new trajectory where $x^{a}\neq x^{a-1}$ for all $a\geq 1$. The update modes in this paper then have the following property: if $x=x^{0}\to\dots\to x^{l}=y$ is a $\mu$-trajectory, then so is its compressed version $x=x^{0}\to\dots\to x^{m}=y$. This property is easy but tedious to prove; as such we shall omit the proof.
For this and the following items, we aim to prove that for a given pair $x,y$ of configurations, there is no trajectory from $x$ to $y$ in a particular update mode (for this item, the most permissive update mode). Based on the considerations above, we always prove that there is no compressed trajectory from $x$ to $y$ in a particular update mode.
Example 3.7 (A trapping reachability pair that is not most permissive).
Let $\mathsf{T}$ be the network from Example 2.3. We claim that $(00,01)$ is a $\mathtt{T}$-reachability pair but not a $\mathtt{M}$-reachability pair. First, we have already given a $\mathtt{T}$-trajectory from $00$ to $01$. Second, any $\mathtt{M}$-trajectory starting at $x^{0}=00$ must have $x^{1}=10$. Therefore, any trajectory reaching $y=01$ must have some $a\geq 2$ with $i^{a}=1$ and $\mathsf{T}_{1}(s^{a})=0$, which is the desired contradiction.
  1. (f)
@@
Example 3.8 (A history-based reachability pair that is not cuttable).
@@
Suppose that there is a trajectory $x=x^{0}\to_{\mathtt{C}}\dots\to_{\mathtt{C}}x^{l}=y$. We start with $x^{0}=000$ with $C^{0}_{i,j}=0$ for all $i,j$. In this network, we must activate the second to activate the third component. For this reason, the first component must be updated. Then, we consider $i^{1}=1$ and $C^{1}=C^{0}$. Accordingly, $s^{1}=x^{0}$ and $x^{1}=100$. Afterwards, to update the second component (i.e., $i^{2}=2$), we must consider $C^{2}_{2,1}=1$ and $C^{2}_{i,j}=C^{1}_{i,j}$ otherwise. As a result, $s^{2}=x^{1}_{1}x^{0}_{2}x^{0}_{3}=100$ and $x^{2}=110$. Later, $x^{3}$ can be calculated with $i^{3}=3$, $C^{3}_{3,2}=2$ and $C^{3}_{i,j}=C^{2}_{i,j}$ otherwise. Accordingly, $s^{3}=x^{1}_{1}x^{2}_{2}x^{0}_{3}=110$ and $x^{3}=111$. At this point, to reach the configuration $101$, we should be able to update the second component by considering a matrix $C^{4}$ where $C^{4}_{2,1}=0$ (to deactivate the second component). However, this is not possible as $C^{4}\geq C^{3}$. In conclusion, $101$ cannot be reached by $000$.
This argument is displayed in the left column of Figure 2.
  1. (g)
@@
Example 3.9 (A cuttable reachability pair that is not history-based).
@@
This argument is displayed in the right column of Figure 2.
  1. (h)
    $\mathtt{I}^{*}\not\subseteq\mathtt{A}^{*}$ and hence $\mathtt{I}\not\subseteq\mathtt{A}$.
Example 3.10 (An interval reachability pair that is not asynchronous).
Let us consider the network $\mathsf{I}$ from Example 2.6. We have shown that $y=011$ is reachable from $x=000$ by interval updates. However, it is easily seen that $011$ is not reachable from $000$ by asynchronous updates.

### v2

@@ 2.2.7 Cuttable updates
The set $E$ is $\{(1,2),(2,3),(1,3)\}$. Let us consider the starting configuration 000 000000\;000 where the first three components are the original ones and the other three correspond to the different elements $E$ in exactly the order presented. If we update the first component, we obtain $f_{1}(\tau^{1}(000\;000))=f_{1}(000)=1$ since $\tau^{1}(000\;000)_{1}=\tau^{1}(000\;000)_{2}=\tau^{1}(000\;000)_{3}=0$. Then, we reach 100 000100\;000. Updating the $(1,2)$ component, we compute 100 100100\;100. Updating now the second one, we have $f_{2}(\tau^{2}(100\;100))=f_{1}(100)=1$ since $\tau^{2}(100\;100)_{1}=1$ and $\tau^{2}(100\;100)_{2}=\tau^{2}(100\;100)_{3}=0$. Then, we reach 110 100110\;100. Updating the $(2,3)$ component, we compute 110 110110\;110. To conclude, we now update the third component, then, $f_{3}(\tau^{3}(110\;110))=f_{1}(010)=1$ since $\tau^{3}(110\;110)_{2}=1$ and $\tau^{3}(110\;110)_{1}=\tau^{3}(110\;110)_{3}=0$. Indeed, we are able to reach 111 110111\;110 given the fact that the third component is able to see the first component at value 0 while the second could see it active.
Cuttable updates $\mathtt{C}$:
All trajectories $(x^{0},\dots,x^{l})$ obtained by $x^{0}\in\mathbb{B}^{n}$ and $w^{a}=(i^{a},s^{a},t^{a})\in[n]\times\mathbb{B}^{n}\times\mathbb{B}^{n}$ for $1\leq a\leq l$, where
  • matrix $C^{a}\in\mathbb{N}^{n\times n}$ such that $C^{0}=0^{n\times n}$, $C^{a}\geq C^{a-1}$ and, for all $i,j\in[n]$, $C^{a}\leq(a-1)^{n\times n}$;
  • source $s^{a}$ with $s^{a}_{j}=x^{C^{a}_{i^{a},j}}_{j}$ for all $j\in[n]$;
  • target $t^{a}=x^{a-1}$.
The correspondence with the initial definition is given in Sect. A.3, and generalizes the one for the interval updates: the memory-based trajectories correspond to the projection over the write nodes.
Example 2.9.
Let us consider the network of the previous example. The asynchronous graph of $\mathsf{C}$ is given here below with an example of a $\mathtt{C}$-trajectory. The same trajectory is also given in Figure 1 in a different format.
000000001001010010011011100100101101110110111111
@@ 0 $\begin{pmatrix}0&0&0\\ 0&0&0\\ 0&0&0\end{pmatrix}$ $000$
11 11 $\begin{pmatrix}0&0&0\\ 0&0&0\\ 0&0&0\end{pmatrix}$ $000$ $000$ $100$
22 22 $\begin{pmatrix}0&0&0\\ 1&0&0\\ 0&0&0\end{pmatrix}$ $100$ $100$ $110$
33 33 $\begin{pmatrix}0&0&0\\ 1&0&0\\ 0&2&0\end{pmatrix}$ $010$ $110$ $111$
We can emphasise how this trajectory corresponds to that of the previous example. Indeed the first component to be activated is the first one that, given the local function, cannot do anything other than activate itself. When updating the second component, the information that the first one has been activated can be transmitted. In the previous example, this is done by updating component $(1,2)$, whereas here it is done by setting the corresponding value of the matrix to $1$ to take into account the first component as an active one. In both cases, this means that the value of the first component causes the second component to activate. When updating the third component, one can decide not to take into account the change of state made by the first component but only consider the second active component. In the previous example, this is done by updating component $(2,3)$ and not updating component $(1,3)$. Here, this corresponds to updating the value of the matrix for the interaction between the second and third components, but not between the first and third. Here too, the result is that the third component changes to $1$.
Figure 1: Possible trajectories of certain BNs according to different update modes (see Examples 2.2 and 2.9).
@@ 3.1 Hierarchy by reachability and trajectory for the different update modes
We are interested in comparing the collections of reachability pairs amongst the different update modes listed above. For any update mode $\mu\in\{\mathtt{A},\mathtt{H},\mathtt{T},\mathtt{M},\mathtt{S},\mathtt{I},\mathtt{C}\}$, we denote its collection of reachability pairs by $\mu^{*}=\{(x,y):x\to_{\mu}^{*}y\}$. Then $\nu$ weakly simulates $\mu$ if $\mu^{*}\subseteq\nu^{*}$ for all BNs. In Figure 2, a strict containment $\mu^{*}\subsetneq\nu^{*}$ means that the containment $\mu^{*}\subseteq\nu^{*}$ holds for all BNs and that there exists a BN where this containment is strict.
Our hierarchy by reachability is based on a hierarchy by trajectory. As such, we also give the full hierarchy of trajectory in Figure 2. Again, a strict containment $\mu\subsetneq\nu$ means that the containment $\mu\subseteq\nu$ holds for all BNs and that there exists a BN where this containment is strict. Clearly, if $\mu\subseteq\nu$, then $\mu^{*}\subseteq\nu^{*}$.
@@
Figure 2: Hierarchy of different update modes. A darker background is used to highlight the update modes introduced for the first time in this work.
The rest of this subsection is devoted to the proof of Theorem 3.1, broken down into items a to i. The structure of the proof is displayed in Figure 3.
$\mathtt{A}^{*}$$\mathtt{I}^{*}$$\mathtt{H}^{*}$$\mathtt{C}^{*}$$\mathtt{M}^{*}$$\mathtt{T}^{*}$$\mathtt{S}^{*}$cdcbcbadfhgi
(a) Hierarchy by reachability
$\mathtt{A}$$\mathtt{H}$$\mathtt{I}$$\mathtt{C}$$\mathtt{M}$$\mathtt{T}$$\mathtt{S}$cabccabaefgh
(b) Hierarchy by trajectory
Figure 3: Structure of the proof of Theorem 3.1. On the edges, the letters identifying the corresponding part of the paper are indicated. Solid blue edges correspond to inclusions, while dotted red edges correspond to non-inclusions.
@@
We now prove that the trapping update mode and the subcube-based update mode allow exactly to reach any configuration in the whole principal trapspace.
@@
Let $x=x^{0}\to_{\mathtt{I}}\dots\to_{\mathtt{I}}x^{l}=y$ be an interval trajectory where $d_{\mathrm{H}}(s^{a-1},s^{a})\leq 1$ for all $1\leq a\leq p-1$ and $d_{\mathrm{H}}(s^{p-1},s^{p})>1$. Let us prove how we can replace the transition $x^{p}\to_{\mathtt{I}}x^{p+1}$ with a sequence of transitions $u^{0}=x^{p}\to_{\mathtt{I}}\dots\to_{\mathtt{I}}u^{n}=x^{p+1}$ and still have an interval trajectory.
@@
is an admissible interval trajectory. The first part, $\dot{x}^{0}\to_{\mathtt{I}}\dots\to_{\mathtt{I}}\dot{x}^{p}$, follows from our hypothesis. For the second part, $\dot{x}^{p}\to_{\mathtt{I}}\dots\to_{\mathtt{I}}\dot{x}^{p+n}$, use $w^{p+i}=(i^{p},v^{i-1},u^{i-1})$. For the third part, $\dot{x}^{p+n}\to_{\mathtt{I}}\dots\to_{\mathtt{I}}\dot{x}^{n+1-l}$, then $\dot{w}^{n-1+a}=w^{a}$ is a valid triple for interval updates. ∎
@@
Let $x=x^{0}\to_{\mathtt{I}}x^{1}\to_{\mathtt{I}}\cdots\to_{\mathtt{I}}x^{l}=y$ be an interval trajectory with corresponding triples $w^{a}=(i^{a},s^{a},t^{a})$ for all $1\leq a\leq l$. According to Lemma 3.4, we assume $d_{\mathrm{H}}(s^{a-1},s^{a})\leq 1$ for all $1\leq a\leq l$.
We first prove, by induction over $1\leq a\leq l$, that there always exists a history-based trajectory $x=x^{0}\to_{\mathtt{H}}s^{1}\to_{\mathtt{H}}\cdots\to_{\mathtt{H}}s^{a}$. The basic case $a=1$ is trivial because $s^{1}=x^{0}$. Assume the statement holds for a1a-1, i.e. $x=x^{0}\to_{\mathtt{H}}s^{1}\to_{\mathtt{H}}\cdots\to_{\mathtt{H}}s^{a-1}$ is a history-based trajectory. Lemma 3.3 settles the case where $s^{a}=s^{a-1}$, hence we assume $s^{a-1}\neq s^{a}$ and, according to Lemma 3.4, $\Delta(s^{a-1},s^{a})=\{j\}$. Then, for $q=V^{a}_{j}$, we have $q<a$ and $x^{q}_{j}=s^{a}_{j}$. Suppose, for the sake of contradiction, that $x^{m}_{j}=x^{q}_{j}$ for all $0\leq m\leq q$. Then, $s^{m}_{j}=x^{q}_{j}$ for all $0\leq m\leq q$, and in particular $s^{a-1}_{j}=s^{a}_{j}$, which is the desired contradiction. Therefore, there exists $q^{\prime}<q$ such that $f_{j}(s^{q^{\prime}})=x^{q}_{j}$. Thus, let $s^{a}=s^{q^{\prime}}$ and $i^{a}=j$. We obtain $(f_{j}(s^{q^{\prime}})=s^{a}_{j},s^{a-1}_{-j})=s^{a}$.
We now prove that we can extend the history-based trajectory all the way to $y$. More formally, we prove that $x=x^{0}\rightarrow s^{1}\rightarrow\cdots\rightarrow s^{l}\rightarrow\cdots\rightarrow y$ is a history-based trajectory, where $s^{l}=g^{l}\rightarrow\cdots\to g^{m}\to\cdots\to g^{l+\mid\Delta\mid}=y$ is any geodesic. Let $\Delta=\Delta(s^{l},y)$ and let $i^{m}$ the coordinate where $g^{m}$ and $g^{m-1}$ differ for all $l+1\leq m\leq l+\mid\Delta\mid$, so that $\{i^{l+1},\ldots,i^{l+\mid\Delta\mid}\}=\Delta$. For all $j\in\Delta$, since $s_{j}^{l}=x^{b}_{j}\neq y_{j}$ (for some $b<l$), there exists $q_{j}$ such that $j=i^{q_{j}}$ and $f_{j}(s^{q_{j}})=y_{j}$. Then $x=x^{0}\rightarrow s^{1}\rightarrow\cdots\rightarrow s^{l}\rightarrow g^{l+1}\to\cdots\rightarrow g^{l+\mid\Delta\mid}=y$ is a history-based trajectory with $(i^{m},s^{m},t^{m})$ for all $l+1\leq m\leq l+\mid\Delta\mid$, where $s^{m}=s^{q_{i^{m}}}$ and $t^{m}=g^{m-1}$. ∎
@@
    $\mathtt{I}\not\subseteq\mathtt{T}$.
Proof (An interval trajectory that is not trapping).
We have already shown that it is an $\mathtt{I}$-trajectory. It is not a $\mathtt{T}$-trajectory, as $x_{1}^{4}\neq x^{3}_{1}$ implies that $i^{4}=1$, while $\mathsf{I}_{1}(x^{0})=\mathsf{I}_{1}(x^{1})=\mathsf{I}_{1}(x^{2})=\mathsf{I}_{1}(x^{3})=1$ implies that $s^{4}\notin\{x^{0},\dots,x^{3}\}$. ∎
  1. (f)
@@
All the update modes considered in this paper have the property that trajectories can be compressed. Our definition of trajectory allows for duplicates, i.e. transitions $x^{a-1}\to_{\mu}x^{a}$ where $x^{a-1}=x^{a}$; a simple example is when $x^{a-1}$ is a fixed point of $f$ and $s^{a}=t^{a}=x^{a-1}$. The update modes in this paper are then closed by removing duplicates.
Formally, let us call a trajectory $y^{0}\to_{\mu}\dots\to_{\mu}y^{l}$ compressed if $y^{a}\neq y^{a-1}$ for all $1\leq a\leq l$. For any trajectory, its compressed version is obtained by removing all duplicates; clearly the latter is indeed compressed.
Lemma 3.6.
For all $\mu\in\{\mathtt{A},\mathtt{H},\mathtt{T},\mathtt{M},\mathtt{S},\mathtt{I},\mathtt{C}\}$, if $x^{0}\to_{\mu}\dots\to_{\mu}x^{l}$ is a $\mu$-trajectory, then so is its compressed version.
Proof.
We only need to prove that one can remove one duplicate of a $\mu$-trajectory and still obtain a $\mu$-trajectory. Let the $\mu$-trajectory $x^{0}\to_{\mu}\dots x^{b-1}\to_{\mu}x^{b}\to_{\mu}x^{l}$ be described by the triples $w^{a}=(i^{a},s^{a},t^{a})$ for all $1\leq a\leq l$, and let the duplicate be $x^{b}=x^{b-1}$. Consider the $\mu$-trajectory $\hat{x}^{0}\to_{\mu}\dots\to_{\mu}\hat{x}^{l-1}$ described by $\hat{w}^{a}=(\hat{i}^{a},\hat{s}^{a},\hat{t}^{a})$, where
[…]
The beginning of the trajectory satisfies $\hat{x}^{0}\to_{\mu}\dots\to_{\mu}\hat{x}^{b-1}=x^{0}\to_{\mu}\dots x^{b-1}$. Then, by induction on $b-1\leq a\leq l-1$, we can easily verify that $\hat{x}^{a}=x^{a-1}$ and that $\hat{x}^{0}\to_{\mu}\dots\to_{\mu}\hat{x}^{a}$ is indeed a $\mu$-trajectory. We conclude that removing the duplicate $x^{b}=x^{b-1}$ still yields a $\mu$-trajectory. ∎
Corollary 3.7.
If there is a $\mu$-trajectory from $x$ to $y$, then there is a compressed $\mu$-trajectory from $x$ to $y$.
For this and the following items, we aim to prove that for a given pair $x,y$ of configurations, there is no $\mu$-trajectory from $x$ to $y$ for a particular update mode $\mu$ (for this item, the most permissive update mode). Based on Corollary 3.7, we always prove that there is no compressed $\mu$-trajectory from $x$ to $y$.
Proof (A trapping reachability pair that is not most permissive).
Let $\mathsf{T}$ be the network from Example 2.3. We claim that $(00,01)$ is a $\mathtt{T}$-reachability pair but not a $\mathtt{M}$-reachability pair. First, we have already given a $\mathtt{T}$-trajectory from $00$ to $01$. Second, any $\mathtt{M}$-trajectory starting at $x^{0}=00$ must have $x^{1}=10$. Therefore, any trajectory reaching $y=01$ must have some $a\geq 2$ with $i^{a}=1$ and $\mathsf{T}_{1}(s^{a})=0$, which is the desired contradiction. ∎
  1. (g)
@@
Proof (A history-based reachability pair that is not cuttable).
@@
Suppose that there is a cuttable trajectory $x=x^{0}\to_{\mathtt{C}}\dots\to_{\mathtt{C}}x^{l}=y$. We start with $x^{0}=000$ with $C^{0}_{i,j}=0$ for all $i,j$. In this network, we must activate the second to activate the third component. For this reason, the first component must be updated. Then, we consider $i^{1}=1$ and $C^{1}=C^{0}$. Accordingly, $s^{1}=x^{0}$ and $x^{1}=100$. Afterwards, to update the second component (i.e., $i^{2}=2$), we must consider $C^{2}_{2,1}=1$ and $C^{2}_{i,j}=C^{1}_{i,j}$ otherwise. As a result, $s^{2}=x^{1}_{1}x^{0}_{2}x^{0}_{3}=100$ and $x^{2}=110$. Later, $x^{3}$ can be calculated with $i^{3}=3$, $C^{3}_{3,2}=2$ and $C^{3}_{i,j}=C^{2}_{i,j}$ otherwise. Accordingly, $s^{3}=x^{1}_{1}x^{2}_{2}x^{0}_{3}=110$ and $x^{3}=111$. At this point, to reach the configuration $101$, we should be able to update the second component by considering a matrix $C^{4}$ where $C^{4}_{2,1}=0$ (to deactivate the second component). However, this is not possible as $C^{4}\geq C^{3}$. In conclusion, $101$ cannot be reached by $000$.
This argument is displayed in the left column of Figure 1. ∎
  1. (h)
@@
Proof (A cuttable reachability pair that is not history-based).
@@
This argument is displayed in the right column of Figure 1. ∎
  1. (i)
    $\mathtt{I}^{*}\not\subseteq\mathtt{A}^{*}$.
Proof (An interval reachability pair that is not asynchronous).
Let us consider the network $\mathsf{I}$ from Example 2.7. We have shown that $y=011$ is reachable from $x=000$ by interval updates. However, it is easily seen that $011$ is not reachable from $000$ by asynchronous updates. ∎